i am three times as old as my niece. seven years ago, i was 4 years less than thrice her age.
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

i am three times as old as my niece. seven years ago, i was 4 years less than thrice her age.
Equations:
I = 3*N
I-7 = 3(N-7)-4
------------------------
Substitute and solve for "N"::
3N-7 = 3N-25
-7 = -25
---------------------------
The equations are contradictory.
No solution.
---
Cheers,
Stan H.
-------------------

Answer by MathTherapy(10552)


i am three times as old as my niece. seven years ago, i was 4 years less than thrice her age.
How can you be 3 times her age now, and some time earlier, 3 times her age, again, less 4?